OpenStudy (anonymous):

someone help me. write two different sequences whose first three terms are 1,2, &4. Describe each pattern. i already have 1,2,4,8,16,32,... ( the number is two times the previous number) i just cant figure out any more and the ones i have thought of i got wrong

OpenStudy (anonymous):

how about 1,2, 4, 7, 11, 16,...

OpenStudy (anonymous):

i tried that answer but my teacher told me it was incorrect

OpenStudy (solomonzelman):

`way 1` \(1,~2,~4,~8,~16,~32\) regular geometric sequence r=2 as you say. `way 2` \(1,~2,~4,~7,~11,~16\) you add +1, +2, +3 and so on as satellite suggested. `way 3` \(1,~2,~4,~9,~23,~64\) \(a_{2}=3a_{1}-1\quad \Longrightarrow \quad 3\cdot 1-1=2\) \(a_{3}=3a_{2}-2\quad \Longrightarrow \quad 3\cdot 2-2=4\) \(a_{4}=3a_{3}-3\quad \Longrightarrow \quad 3\cdot 4-3=9\) \(a_{5}=3a_{4}-4\quad \Longrightarrow \quad 3\cdot 9-4=23\) \(a_{6}=3a_{5}-5\quad \Longrightarrow \quad 3\cdot 23-5=64\) this satisfies the formula, \(a_{n+1}=3a_{n}-n\)

OpenStudy (solomonzelman):

With only three numbers you can make almost any pattern...

OpenStudy (anonymous):

thank you